# Lightening the Load: A Cluster-Based Framework for A Lower-Overhead, Provable Website Fingerprinting Defense

### v1

@@ Abstract
Website fingerprinting (WF) attacks remain a significant threat to encrypted traffic, prompting the development of a wide range of defenses. Among these, two prominent classes are regularization-based defenses, which shape traffic using fixed padding rules, and supersequence-based approaches, which conceal traces among predefined patterns. In this work, we present a unified framework for designing an adaptive WF defense that combines the effectiveness of regularization with the provable security of supersequence-style grouping. The scheme first extracts behavioural patterns from traces and clusters them into $(k,l)$-diverse anonymity sets; an early-time-series classifier (adapted from ECDIRE) then switches from a conservative global set of regularization parameters to the lighter, set-specific parameters. We instantiate the design as Adaptive Tamaraw, a variant of Tamaraw that assigns padding parameters on a per-cluster basis while retaining its original information-theoretic guarantee. Comprehensive experiments on public real-world datasets confirm the benefits. By tuning $k$, operators can trade privacy for efficiency: in its high-privacy mode Adaptive Tamaraw pushes the bound on any attacker’s accuracy below 30%, whereas in efficiency-centred settings it cuts total overhead by 99% compared with classic Tamaraw.
@@ I Introduction
  • Our experiments show that Adaptive Tamaraw offers flexible tunability, enabling practitioners to adjust the defense parameters to meet specific privacy and efficiency requirements. When configured for strong privacy, the defense reduces attack accuracy to well below 30%. In efficiency-focused settings, it achieves substantial reductions in the total overhead, up to 99% relative to the original Tamaraw. The implementation is publicly available at https://anonymous.4open.science/r/Adaptive-Tamaraw-F864.
@@ II-B1 Empirical WF Defenses
A large body of website fingerprinting defenses rely on empirical evaluation to demonstrate effectiveness against specific attacks. These defenses often succeed in reducing fingerprinting accuracy and can be tuned for lower overhead. However, they typically lack formal guarantees and remain vulnerable to stronger or more adaptive adversaries. Empirical defenses can be broadly categorized into two main families:
@@ II-B2 WF Defenses With Formal Bounds
As summarized in Table I, defenses offering formal bounds achieve low attack accuracies even after many years. Adaptive Tamaraw preserves provable guarantees while introducing clustering and adaptive switching mechanisms to reduce overhead. Crucially, unlike supersequence methods, it remains applicable to websites beyond the training set, making it suitable for real-world, out-of-training scenarios.
@@ III Threat Model
Following prior work [15, 14, 22], we assume a local, passive adversary who can observe all traffic between the client and the guard node in Tor. The adversary cannot alter traffic but can record packet size, timing, and direction to infer the visited webpage. We also assume the adversary is capable of segmenting traffic into individual page loads, and users visit one page at a time. This assumption is standard in WF evaluations to simplify analysis and avoid assumptions that favor the defense.
To carry out the attack, the adversary collects training data by visiting a set of monitored webpages (i.e., those of particular interest) and recording their traffic traces under conditions similar to the target user’s environment. The adversary’s goal is to recognize which of these monitored sites the client is loading.
@@ IV Problem Statement
A defense modifies the trace $f$ to produce a defended trace $f^{\prime}$ by delaying real packets ($t_{i}\mapsto t_{i}+\theta_{i}$) or adding dummies $(t^{\prime},d^{\prime})$, obscuring patterns used by WF attacks.
We evaluate defenses using:
  • Bandwidth Overhead:
    $B(f,f^{\prime})=\frac{|f^{\prime}|-N}{N},$
    the ratio of dummy to real packets.
  • Time Overhead:
    $T(f,f^{\prime})=\frac{t^{\prime}-t_{N}}{t_{N}},$
    the relative delay of the last real packet, where $t^{\prime}$ is the timestamp of the last real packet in the defended trace.
@@ V-C Intra-Webpage Pattern Detection
With traces represented as TAM time series, we cluster the traces of each webpage individually to isolate recurring traffic patterns. We refer to the resulting clusters as intra-webpage clusters (i.e., for the remainder of this section, when we refer to intra-webpage clusters, we mean clusters obtained by grouping traces originating from a single webpage). To perform this task, we adopt the Cluster Affinity Search Technique (CAST) [26]. CAST does not require us to specify the number of clusters in advance; instead, it uses an affinity threshold $T$ (which is a hyperparameter) to determine membership. The affinity of a time series $F_{x}$ with respect to a particular intra-webpage cluster $\mathcal{S}$ is defined as follows:
$a_{\mathcal{S}}(F_{x})\;=\;\frac{1}{\lvert\mathcal{S}\rvert}\sum_{F_{y}\,\in\,\mathcal{S}}A(F_{x},F_{y}).$
where $A(F_{x},F_{y})$ is the similarity between time series $F_{x}$ and $F_{y}$, and $\lvert\mathcal{S}\rvert$ is the size of the intra-webpage cluster. CAST evolves a cluster by repeatedly selecting the unassigned time series with the highest affinity to the cluster:
  • If its affinity exceeds the fixed threshold $T$, the series is added to the intra-webpage cluster.
  • Otherwise, the algorithm checks if any existing member in the intra-webpage cluster has an affinity falling below the threshold; if so, that low-affinity member is removed.
  • If no member meets the removal criterion, the intra-webpage cluster is deemed complete (closed), and CAST proceeds to form a new intra-webpage cluster with the remaining data.
Our preliminary evaluations revealed that for traces in a single webpage, CAST generated several large intra-webpage clusters, along with a long tail of smaller ones. Small intra-webpage clusters are not suitable, as our goal is to regularize all traces within an intra-webpage cluster to appear identical; if an intra-webpage cluster contains too few traces, the defense incurs the overhead of padding and shaping with minimal privacy benefit. To be more precise, the attacker’s uncertainty remains low despite the added cost, resulting in an inefficient overhead-to-privacy tradeoff. To produce better clusters, we introduced four modifications to CAST, which we will describe below.
@@
We apply the modified CAST algorithm separately to each webpage in the dataset to extract a set of characteristic traffic patterns. These extracted patterns form the foundation for all subsequent phases. The complete pseudocode for this phase is given in Appendix D.
@@ V-D Anonymity Set Generation
This formulation implicitly enhances $l$-diversity, since adding a pattern from a new webpage typically reduces the attacker’s success rate based on the webpage in the majority, thus lowering $d(C,p)$. We provide empirical validations for this attribute in Appendix B-C. During clustering, we therefore always merge the candidate $p$ that minimizes $d(C,p)$, encouraging both homogeneity in regularized shape and diversity in origin. Algorithm 1 provides detailed pseudocode of our proposed anonymity set generation algorithm.
@@ V-F Security Bound for Adaptive Tamaraw
Despite transitioning from a fixed global defense to lighter, AS-specific configurations after a safe prediction point, Adaptive Tamaraw retains a provable security guarantee. Specifically, we show that the defense remains non-uniformly weighted $\delta$-non-injective, ensuring that the attacker’s average success probability is formally bounded.
@@
The formal proof of this result is presented in Appendix C.
This result is intuitive: the attacker’s overall bounded success rate is the average across the bound for each combination of anonymity set and trace length. Any two traces from the same set and of the same length are made indistinguishable by the defense. Consequently, the attacker’s optimal strategy in an anonymity set reduces to guessing the most common webpage in each such length-matched group, which limits their average success rate captured by $\bar{A}(S_{i})$.
@@ VI Experimental Evaluation
In this section, we empirically evaluate the effectiveness, efficiency, and generalizability of Adaptive Tamaraw. Our goal is to demonstrate that the proposed defense reduces bandwidth and latency overhead while maintaining provable bounds on attacker success rates, even under strong adversarial models. Section VI-A describes the experimental setup, including datasets, hyperparameters, and model architectures. Section VI-B evaluates performance on in-training webpages, where all defended sites were included in the training phase. Section VI-C extends the analysis to out-of-training webpages, assessing the defense’s robustness against previously unseen traffic patterns. Finally, Section VI-D verifies the tightness of the theoretical bounds established in Section V-F, comparing them against observed attack accuracies.
@@ VI-A Experimental Setup
We conduct our experiments using the widely adopted Tor-based website fingerprinting dataset originally collected by [6], which has become a standard benchmark in recent WF research, including works such as [8, 22, 41, 42]. The dataset comprises two parts: a closed-world set and an open-world set. The closed-world subset includes 1,000 traffic traces for each of 95 monitored webpages, totaling 95,000 labeled instances. All traces were collected via real Tor connections and reflect authentic network conditions including congestion, jitter, and Tor circuit variability.
@@
For each $L$, we explore a grid of candidate values for $\rho_{\text{in}}$ and $\rho_{\text{out}}$. Given that incoming traffic typically arrives more frequently than outgoing traffic during webpage loading [10, 12], we sweep $\rho_{\text{in}}$ in the range $[0.001,0.006]$ and $\rho_{\text{out}}$ in the range $[0.005,0.21]$. We retain only the Pareto-optimal configurations (those for which no other configuration achieves strictly lower bandwidth and time overhead simultaneously [10]). This selection yields 33 unique $(\rho_{\text{in}},\rho_{\text{out}})$ pairs that span the optimal trade-off frontier between efficiency and security. A detailed report of these configurations is provided in Appendix B-A.
@@
To ensure optimal performance and fair evaluation, we conducted a grid search for each hyperparameter over its respective search space. Table II summarizes the key hyperparameters used in Adaptive Tamaraw and its underlying components. The first three parameters ($k$, Max Intra-Webpage Patterns, and $\alpha$) were selected to minimize overhead while preserving security guarantees. The remaining model-specific hyperparameters for the Holmes and kFP classifiers were chosen to maximize prediction accuracy in the early anonymity set detection phase.
@@ VI-B In-Training Evaluation
In the in-training setting, we evaluate the effectiveness of our defense in protecting webpages that were explicitly included in the defense’s training dataset. Specifically, we focus on the 95 monitored webpages from the closed-world portion of the dataset introduced in Section VI-A. These sites serve as our set of protected targets. Following prior work [22, 8], we partition the dataset into training, validation, and testing sets using an 8:1:1 ratio. At a high level, our full defense pipeline proceeds as follows. We begin by applying the modified CAST algorithm V-C to the training traces of each webpage to extract recurring traffic patterns. These patterns are then clustered using the $k$-anonymity-based algorithm introduced in Section V-D to obtain our anonimity-sets (we perform the experiments by varying the minimum required anonymity set size $k$ from 2 to 30). For each of the obtained anonymity-sets, we identify a safe timestamp using early time-series classification as described in Section V-E, allowing us to switch from a global Tamaraw configuration to a cluster-specific configuration as early as possible. Finally, we compare the performance against applying the classical Tamaraw to the traces.
@@
We proceed to evaluate the performance of Adaptive Tamaraw using the held-out test set. For each test trace $i$, and for each global Pareto regularization pair $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$, we first apply Tamaraw with the global parameters and record the resulting bandwidth and time overheads. Next, we evaluate the same trace under the complete adaptive Tamaraw strategy: we begin with $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$ and switch to the specific configuration of the anonymity set $(\rho_{\text{in}}^{\mathcal{S}_{i}},\rho_{\text{out}}^{\mathcal{S}_{i}})$ once the predicted anonymity set becomes identifiable according to its safe prediction time. We then record the resulting overheads. We repeat this procedure for all the 33 global Pareto regularization pairs $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$.
Table III presents average total overhead (combined time and bandwidth, averaged over all global regularization pairs) for static Tamaraw versus Adaptive Tamaraw at various cluster sizes $k$, across three bucket lengths $L\in\{100,500,1000\}$. Adaptive Tamaraw achieves notable reductions across all settings, particularly for small $k$. For example, when L=1000L{=}1000, Adaptive Tamaraw reduces the average overhead from 459% (Tamaraw) to 360% at k=2k{=}2, a drop of 99%. Even at k=30k{=}30, the reduction remains significant at 60%. For smaller $L$, the relative savings are smaller but still meaningful: at L=100L{=}100 and k=2k{=}2, the overhead drops from 243% to 205% (38% decrease), while at L=500L{=}500, the drop is from 299% to 258% (41%).
The choice of $k$ controls the defense’s effectiveness: smaller values result in a looser bound on attacker accuracy (fully explored in Section VI-D) while producing the smallest overhead. Interestingly, Adaptive Tamaraw results in higher overhead savings with smaller $k$, i.e. when the anonimity sets cover less intra-webpage patterns. An intuitive explanation is that in high $k$ settings, each anonymity set is large and therefore does not benefit as much from localized, set-specific adaptive parameters. More comprehensive results across a broader range of $k$ values are provided in Appendix B-B.
TABLE III: Average total overhead (time + bandwidth) comparison between static Tamaraw and Adaptive Tamaraw across multiple cluster sizes $k$ and bucket lengths $L$. Each entry shows the absolute overhead percentage, with the difference from Tamaraw in parentheses. Adaptive Tamaraw consistently reduces overhead, especially for small values of $k$, while maintaining security guarantees.
L Tamaraw (Baseline) Adaptive Tamaraw
@@
100 243.0% 205.0% (-38.0) 217.0% (-26.0) 235.0% (-8.0)
500 299.0% 258.0% (-41.0) 270.0% (-29.0) 283.0% (-16.0)
1000 459.0% 360.0% (-99.0) 379.0% (-80.0) 399.0% (-60.0)
Per-Trace Analysis. Beyond aggregate averages, we examine how Adaptive Tamaraw impacts overhead on a per-trace basis. To this end, we selected a representative global Pareto configuration, $(\rho_{\text{in}}=0.012,\rho_{\text{out}}=0.04)$, which was also reported in the original Tamaraw paper [10], and offers a balanced trade-off between bandwidth and time overhead. We applied both static Tamaraw and Adaptive Tamaraw (with $k=7$ and $L=100$) to each test trace using this configuration as the global parameters. The resulting average bandwidth and time overheads for static Tamaraw were 106% and 53%, respectively. With Adaptive Tamaraw, these dropped to 95% and 47%, corresponding to an overall overhead reduction of approximately 17%. We measured the difference in total overhead between the two methods for each trace and plotted the distribution.

### v2

@@ Abstract
Website fingerprinting (WF) attacks remain a significant threat to encrypted traffic, prompting the development of a wide range of defenses. Among these, two prominent classes are regularization-based defenses, which shape traffic using fixed padding rules, and supersequence-based approaches, which conceal traces among predefined patterns. In this work, we present a unified framework for designing an adaptive WF defense that combines the effectiveness of regularization with the provable security of supersequence-style grouping. The scheme first extracts behavioural patterns from traces and clusters them into $(k,l)$-diverse anonymity sets; an early-time-series classifier (adapted from ECDIRE) then switches from a conservative global set of regularization parameters to the lighter, set-specific parameters. We instantiate the design as Adaptive Tamaraw, a variant of Tamaraw that assigns padding parameters on a per-cluster basis while retaining its original information-theoretic guarantee. Comprehensive experiments on public real-world datasets confirm the benefits. By tuning $k$, operators can trade privacy for efficiency: in its high-privacy mode, Adaptive Tamaraw pushes the bound on any attacker’s accuracy below 30%, whereas in efficiency-centred settings it cuts total overhead by 99 percentage points compared with classic Tamaraw.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.241760 www.ndss-symposium.org
@@ I Introduction
  • Our experiments show that Adaptive Tamaraw offers flexible tunability, enabling practitioners to adjust the defense parameters to meet privacy and efficiency requirements. When configured for strong privacy, the defense reduces attack accuracy to below 30%. In efficiency-focused settings, it achieves substantial reductions in the total overhead, up to 99 percentage points relative to the original Tamaraw. The implementation is publicly available at https://github.com/khashayarkhaj/Adaptive-Tamaraw.
@@ II-B1 Empirical WF Defenses
Many website fingerprinting defenses rely on empirical evaluation to show effectiveness against specific attacks. They often reduce fingerprinting accuracy and can be tuned for lower overhead but lack formal guarantees and remain vulnerable to stronger or adaptive adversaries. Empirical defenses fall into two main categories:
@@ II-B2 WF Defenses With Formal Bounds
As summarized in Table I, a clear dichotomy exists in the literature: while many low-overhead defenses have been proposed, those lacking formal guarantees have often been broken by subsequent attacks. In contrast, defenses with formal bounds, such as Tamaraw, have proven resilient over time but at the cost of high overhead. The core novelty of our paper lies in introducing a framework to bridge this gap between provable security and practical efficiency. In this work, we build upon Tamaraw specifically due to its unique and unbroken provably secure foundation, and demonstrate that its overhead can be reduced via dynamic clustering and switching mechanisms, all while preserving the original formal guarantees. Crucially, unlike supersequence methods, this approach remains applicable to websites beyond the training set, making it suitable for real-world, out-of-training scenarios.
@@ III Threat Model
Following prior work [14, 13, 34], we assume a local, passive adversary who observes all traffic between the client and the Tor guard node. The adversary cannot modify traffic but records packet size, timing, and direction to infer the visited webpage. They can segment traffic into individual page loads, assuming users visit one page at a time. To execute the attack, the adversary collects training data by visiting monitored webpages and recording their traffic under similar conditions to the target user. The goal is to identify which monitored site the client is loading.
@@
We also assume the defense operates without prior knowledge of the destination webpage. Our design goal is to develop an easy-to-implement defense that is generally suitable for browsing any site, including dynamic webpages generated on the fly. As such, we do not limit the defense’s applicability to a pre-configured set of static pages and we also avoid the privacy risk of disclosing the destination to a defending proxy. As we will elaborate upon in Section V-C, our approach is designed to operate on the more fundamental and recurring traffic patterns that emerge during a page load.
@@ IV Problem Statement
A defense modifies the trace $f$ to produce a defended trace $f^{\prime}$ by delaying real packets or adding dummies, obscuring patterns used by WF attacks. We evaluate the overhead of defenses by bandwidth overhead and time overhead. The bandwidth overhead is measured as the total number of dummy packets divided by the total number of real packets over the whole dataset. Similarly, the time overhead is measured as the total extra time divided by the total loading time in the undefended case over the whole dataset.
@@ V-C Intra-Webpage Pattern Detection
With traces represented as TAM time series, we cluster the traces of each webpage individually to isolate recurring traffic patterns. We refer to the resulting clusters as intra-webpage clusters (i.e., for the remainder of this section, when we refer to intra-webpage clusters, we mean clusters obtained by grouping traces originating from a single webpage).
To perform this clustering, we adopt the Cluster Affinity Search Technique (CAST) [1]. CAST is a similarity-based clustering algorithm that avoids pre-specifying the number of clusters. Instead, it incrementally builds clusters by comparing each trace’s affinity (its average similarity) to the current cluster. We define the affinity of a time series $F_{x}$ to a cluster $S$ as:
$a_{S}(F_{x})=\frac{1}{|S|}\sum_{F_{y}\in S}A(F_{x},F_{y}),$ (2)
where $A(F_{x},F_{y})$ is the similarity measure between two TAM series. Intuitively, $a_{S}(F_{x})$ captures how well $F_{x}$ fits with the existing members of $S$: high affinity indicates that $F_{x}$ behaves similarly to most traces in $S$, while low affinity indicates that it is not consistent with the dominant traffic pattern in that cluster. CAST maintains a currently “open” cluster and evolves it using a fixed affinity threshold $T$. At each step, the unassigned trace with the highest affinity to the open cluster is examined:
  • Addition: If its affinity is at least $T$, the trace is added to the cluster.
  • Removal: Otherwise, CAST checks if any current member of the cluster has affinity below $T$ with the current cluster. Such traces are considered weakly related to the cluster and are removed.
  • Closure: If no addition or removal is possible, the cluster is considered stable and is closed; CAST then begins forming the next cluster from the remaining traces.
Our preliminary evaluations (discussed in Appendix C-B) revealed that for traces in a single webpage, CAST generated several large intra-webpage clusters, along with a long tail of smaller ones. Small intra-webpage clusters are not suitable, as our goal is to regularize all traces within an intra-webpage cluster to appear identical; if an intra-webpage cluster contains too few traces, the defense incurs the overhead of padding and shaping with minimal privacy benefit. To be more precise, the attacker’s uncertainty remains low despite the added cost, resulting in an inefficient overhead-to-privacy tradeoff. To produce better clusters, we introduced four modifications to the original CAST algorithm proposed in [1], which we will describe below.
@@
We apply the modified CAST algorithm separately to each webpage in the dataset to extract a set of characteristic traffic patterns. These extracted patterns form the foundation for all subsequent phases.
@@ V-D Anonymity Set Generation
This formulation implicitly enhances $l$-diversity, since adding a pattern from a new webpage typically reduces the attacker’s success rate based on the webpage in the majority, thus lowering $d(C,p)$. We provide empirical validations for this attribute in Appendix C-A. During clustering, we therefore always merge the candidate $p$ that minimizes $d(C,p)$, encouraging both homogeneity in regularized shape and diversity in origin. Algorithm 1 provides detailed pseudocode of our proposed anonymity set generation algorithm.
@@ V-F Security Bound for Adaptive Tamaraw
Despite transitioning from a fixed global defense to lighter, AS-specific configurations after a safe prediction point, Adaptive Tamaraw retains a provable security guarantee. Specifically, we show that the defense remains non-uniformly weighted $\delta$-non-injective, ensuring that the attacker’s average success probability is formally bounded. This bound reflects the attacker’s expected success probability across the entire distribution of defended traffic, aligned with standard metrics from prior work, including Tamaraw  [6], which included non-uniform (i.e., average) security.
@@
The formal proof of this result is presented in Appendix E.
This result is intuitive: the attacker’s overall bounded success rate is the average across the bound for each combination of anonymity set and trace length. Any two traces from the same set and of the same length are made indistinguishable by the defense, a principle inherited from Tamaraw. Consequently, the attacker’s optimal strategy in an anonymity set reduces to guessing the most common webpage in each such length-matched group, which limits their average success rate captured by $\bar{A}(S_{i})$. Our novelty lies in deriving the overall security bound by averaging these individual success probabilities across all possible anonymity sets.
@@ VI Experimental Evaluation
This section empirically evaluates the effectiveness, efficiency, and generalizability of Adaptive Tamaraw. The goal is to show that the defense lowers bandwidth and latency overhead while maintaining provable bounds on attacker success, even under strong adversaries. Section VI-A outlines the experimental setup; Section VI-B tests performance on in-training webpages (defended sites seen during training); Section VI-C examines robustness on out-of-training webpages with unseen traffic; and Section VI-D checks the tightness of theoretical bounds from Section V-F against observed attack accuracies.
@@ VI-A Experimental Setup
We conduct our experiments on two well-known public WF datasets. The first is collected by Sirinam et al. [36], and it has become a standard benchmark in recent WF research [33, 34, 46, 9]. For our experiments, we used its closed-world subset, which contains 1,000 traffic traces for each of the 95 monitored websites. The second is the large-scale Automated Website Fingerprinting (AWF) dataset from Rimmer et al. [31]. From this collection, we use the subset corresponding to the top 100 most popular websites, which provides 2,500 traffic traces per website. Specifically, both data sets consist of traces collected over live Tor connections, reflecting authentic and variable network conditions, including congestion, jitter, and circuit diversity.
@@
For each $L$, we explore a grid of candidate values for $\rho_{\text{in}}$ and $\rho_{\text{out}}$. Given that incoming traffic typically arrives more frequently than outgoing traffic during webpage loading [6, 18], we sweep $\rho_{\text{in}}$ in the range $[0.001,0.006]$ and $\rho_{\text{out}}$ in the range $[0.005,0.21]$. We retain only the Pareto-optimal configurations (those for which no other configuration achieves strictly lower bandwidth and time overhead simultaneously [6]). This selection yields 33 unique $(\rho_{\text{in}},\rho_{\text{out}})$ pairs for the Sirinam et al. dataset and 40 for the AWF dataset, that span the optimal trade-off frontier between efficiency and security. A detailed report of these configurations is provided in Appendix B-A.
@@
For optimal performance and fair evaluation, we performed a grid search over each hyperparameter. Table II lists the key parameters of Adaptive Tamaraw. The first three ($k$, Max Intra-Webpage Patterns, and $\alpha$) minimize overhead while maintaining security, and the others maximize prediction accuracy for early anonymity set detection.
@@ VI-B In-Training Evaluation
In the in-training setting, we evaluate the effectiveness of our defense in protecting webpages that were explicitly included in the defense’s training dataset. Specifically, we conduct our evaluation on the monitored websites from both the Sirinam et al. and AWF datasets. These sites serve as our set of protected targets. Following prior work [34, 33], we partition the dataset into training, validation, and testing sets using an 8:1:1 ratio. At a high level, our full defense pipeline proceeds as follows. We begin by applying the modified CAST algorithm V-C to the training traces of each webpage to extract recurring traffic patterns. These patterns are then clustered using the $k$-anonymity-based algorithm introduced in Section V-D to obtain our anonimity-sets (we perform the experiments by varying the minimum required anonymity set size $k$ from 2 to 30). For each of the obtained anonymity-sets, we identify a safe timestamp using early time-series classification as described in Section V-E, allowing us to switch from a global Tamaraw configuration to a cluster-specific configuration as early as possible. Finally, we compare the performance against applying the classical Tamaraw to the traces.
@@
We proceed to evaluate the performance of Adaptive Tamaraw using the held-out test set. For each test trace $i$, and for each global Pareto regularization pair $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$, we first apply Tamaraw with the global parameters and record the resulting bandwidth and time overheads. Next, we evaluate the same trace under the complete adaptive Tamaraw strategy: we begin with $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$ and switch to the specific configuration of the anonymity set $(\rho_{\text{in}}^{\mathcal{S}_{i}},\rho_{\text{out}}^{\mathcal{S}_{i}})$ once the predicted anonymity set becomes identifiable according to its safe prediction time. We then record the resulting overheads. We repeat this procedure for all the global Pareto regularization pairs $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$.
Table III presents the average overheads (time and bandwidth, each averaged over all global regularization pairs) for static Tamaraw versus Adaptive Tamaraw at various cluster sizes $k$, across three bucket lengths $L\in\{100,500,1000\}$ for both datasets.. On the Sirinam et al. dataset, the savings are substantial: when $L=1000$, Adaptive Tamaraw reduces the average overheads from 258% bandwidth and 199% time overhead to 223% and 135% respectively at $k=2$, a combined reduction of 99.0 percentage points. For the AWF dataset, we observe that while the baseline overhead for static Tamaraw is generally higher, the overhead reductions from our defense are still pronounced. For instance, at $L=1000$ and $k=2$ on AWF, the average overheads drop from 182% and 207% to 145% and 154%, a reduction of 90.0 percentage points.
The choice of $k$ controls the defense’s effectiveness: smaller values result in a looser bound on attacker accuracy (fully explored in Section VI-D) while producing the smallest overhead. Interestingly, Adaptive Tamaraw results in higher overhead savings with smaller $k$, i.e. when the anonymity sets cover fewer intra-webpage patterns. An intuitive explanation is that in high $k$ settings, each anonymity set is large and therefore does not benefit as much from localized, set-specific adaptive parameters. More comprehensive results across a broader range of $k$ values are provided in Appendix B-D.
TABLE III: Average overheads (time and bandwidth) comparison between static Tamaraw and Adaptive Tamaraw across multiple cluster sizes $k$ and bucket lengths $L$. Each entry shows the absolute overhead percentage, with the difference from Tamaraw in parentheses. Adaptive Tamaraw consistently reduces overhead, especially for small values of $k$.
Dataset L Tamaraw (Baseline) Adaptive Tamaraw
@@
BW Time BW Time BW Time BW Time
Sirinam et al. 100 158% 83% 136% (-22) 68% (-15) 144% (-14) 74% (-9) 157% (1) 77% (-6)
500 198% 98% 176% (-22) 86% (-12) 184% (-14) 86% (-11) 196% (-2) 87% (-11)
1000 258% 199% 223% (-35) 135% (-64) 235% (-23) 144% (-55) 248% (-10) 150% (-49)
AWF 100 151% 153% 100% (-51) 111% (-42) 109% (-42) 123% (-30) 125% (-26) 135% (-18)
500 157% 183% 122% (-35) 147% (-36) 132% (-25) 155% (-28) 141% (-16) 172% (-6)
1000 182% 207% 145% (-37) 154% (-53) 157% (-25) 162% (-45) 169% (-13) 182% (-25)
Per-Trace Analysis. Beyond aggregate averages, we examine how Adaptive Tamaraw impacts overhead on a per-trace basis. To this end, we selected a representative global Pareto configuration, $(\rho_{\text{in}}=0.012,\rho_{\text{out}}=0.04)$, which was also reported in the original Tamaraw paper [6], and offers a balanced trade-off between bandwidth and time overhead. We applied both static Tamaraw and Adaptive Tamaraw (with $k=7$ and $L=100$) to each test trace in the Sirinam et al. dataset using this configuration as the global parameters. The resulting average bandwidth and time overheads for static Tamaraw were 106% and 53%, respectively. With Adaptive Tamaraw, these dropped to 95% and 47%, corresponding to an overall overhead reduction of approximately 17%. We measured the difference in total overhead between the two methods for each trace and plotted the distribution.